what is the sum of the first 44 positive odd numbers
Found 2 solutions by Alan3354, Boreal:
Answer by Alan3354(69443)

what is the sum of the first 44 positive odd numbers
-----
44^2 = 1936

Answer by Boreal(15235)

The sum of the first n odd numbers is n^2
This can be proven, but for 5 odd numbers, 1,3,5,7,9, the sum should be 5^2, or 25, which it is
So for the first 44 odd numbers, the sum is n^2=44^2=1936